2. Using the formulae for geometric progression

(a) The value of a lathe originally valued at £9000 depreciates by 5% each year. Calculate its value after 6 years.

(b) The lathe will be sold when its value falls below £6000 after how many years will this happen?


3) A construction company is fined for each day of delay in the construction of a

bridge. The penalty for the delay is £3000 for the first day and then an additional

£500 for each subsequent day ( i.e. Day 1 £3000, day 2 £3500, day 3 £4000

etc.). The maximum fine the company can afford is £75000.

Use arithmetic sequences formulae to find the maximum number of days they can

afford to delay.


4) Simplify the following, using the rectangular form of complex numbers

5 + 2j / 7 - 3j  and (5+ 3j) (-8-2j)


5) Simplify, using the polar form of complex numbers

a. 5 < 55 ^degree x 8 < - 28^ degree


b. 5 < 55 ^ degree / 8 < - 28 ^ degree

Convert both answers to rectangular form



1
Expert's answer
2021-05-03T05:39:29-0400

2. (a)


"A_n=A_0(1-r)^n"

"A_6=£9000\cdot(1-0.05)^6=£6615.83"

(b)

"A_n=A_0(1-r)^n"

"9000(1-0.05)^n<6000"

"0.95^n<\dfrac{2}{3}"

"n>\dfrac{\ln(\dfrac{2}{3})}{\ln0.95}"

"n=8"

"\u00a39000(1-0.05)^7=\u00a36285.04"


"\u00a39000(1-0.05)^8=\u00a35770.78"


The lathe will be sold after 8 years.


3.


"a_1=\u00a33000, d=\u00a3500"

"a_n=a_1+d(n-1)""S_n=n\cdot\dfrac{2a_1+d(n-1)}{2}"

Given "S_n=\u00a375000"


"n\cdot\dfrac{2(3000)+500(n-1)}{2}\leq75000"

"12n+n^2-n\leq300"

"n^2+11n-300\leq0"

"n=\dfrac{-11\pm\sqrt{1321}}{2}, n>0"

"\dfrac{-11\pm\sqrt{1321}}{12}\approx12.673"


"n=12"


The maximum number of days they can afford to delay is 12.


4)

"\dfrac{5+2j}{7-3j}=\dfrac{(5+2j)(7+3j)}{7^2+3^2}=\dfrac{29+29j}{58}"

"=\dfrac{1}{2}+\dfrac{1}{2}j"


"(5+3j)(-8-2j)=-40-10j-24j+6"

"=-34-34j"

5) Simplify, using the polar form of complex numbers


"5e^{j55\degree}\times8e^{-j28\degree}=40e^{j27\degree}"

"=40\cos(27\degree)+j40\sin(27\degree)"

b.

"\dfrac{5e^{j55\degree}}{8e^{-j28\degree}}=\dfrac{5}{8}e^{j83\degree}"


"=\dfrac{5}{8}\cos(83\degree)+j\dfrac{5}{8}\sin(83\degree)"